# Evaluating Creative Short Story Generation in Humans and Large Language Models

### v4

@@ Abstract
Story-writing is a fundamental aspect of human imagination, relying heavily on creativity to produce narratives that are novel, effective, and surprising. While large language models (LLMs) have demonstrated the ability to generate high-quality stories, their creative story-writing capabilities remain under-explored. In this work, we conduct a systematic analysis of creativity in short story generation across 60 LLMs and 60 people using a five-sentence creative story-writing task. We use measures to automatically evaluate model- and human-generated stories across several dimensions of creativity, including novelty, surprise, diversity, and linguistic complexity. We also collect creativity ratings and Turing Test classifications from non-expert and expert human raters and LLMs. Automated metrics show that LLMs generate stylistically complex stories, but tend to fall short in terms of novelty, surprise and diversity when compared to average human writers. Expert ratings generally coincide with automated metrics. However, LLMs and non-experts rate LLM stories to be more creative than human-generated stories. We discuss why and how these differences in ratings occur, and their implications for both human and artificial creativity.
@@ Introduction
Story-writing lies at the core of human imagination and communication, serving as a potent means to connect and convey ideas effectively spanning across all human cultures and time periods (?). It typically demands creativity, especially when shaping a captivating and persuasive narrative. Creativity is the ability to produce novel, useful, and surprising ideas, and has been widely studied as a crucial aspect of human cognition (????). While humans are natural storytellers, getting machines to generate stories automatically has been a long-time challenge (?????). However, recently large language models (?) have been shown to produce high-quality short and long stories on arbitrary topics (??). These stories are often evaluated by humans on their long-term coherence, relevance to the premise, repetitiveness, and general interestingness to the reader (?). However, the extent to which these model-generated stories are truly creative, i.e. novel, effective, and surprising, remains under-studied.
@@
Recent works evaluating LLM’s ability to produce creative content have shown that models largely fall behind professional human writers (????). On the other hand, Orwig et al. (?) finds no significant difference between average human and AI-generated short stories in terms of creativity ratings by non-experts or GPT-4, when comparing humans to ChatGPT models. The extent to which these results hold when considering a collection of different models and evaluations across multiple dimensions of creativity remains unclear.
To bridge these gaps, in this work, we conduct a systematic analysis of creativity in short story generation in humans and LLMs. We employ a creative short story generation task that is typically used in psychology to measure the creativity of humans (???). In this task, the goal is to write a short creative story in approximately five sentences based on three cue words such as stamp, letter and send. We evaluate 60 humans and 60 state-of-the-art instruction-finetuned large language models on this task and analyze their performance based on multiple dimensions of creativity representing, overall creativity, and the specific aspects of novelty, surprise, diversity, and complexity. Our analysis shows that model-generated stories tend to employ linguistically more complex structures than humans; however, they significantly fall short when it comes to novelty, diversity, and surprise compared to average human writers.
@@
Evaluating creativity is a challenging task due to its subjective nature, however, several evaluation methods have been proposed in the past (??). The most common method of creativity evaluation is called the Consensual Assessment Technique (CAT) (?). CAT relies on the collective judgment of human experts. However, the level of expertise required for a human rater is a subject of debate, with most evidence favoring experts over non-experts as good raters (???????). With the rise of powerful generative AI models, LLMs are increasingly being used as judges in many evaluation tasks including creativity (??????). Consequently, in addition to automated metrics, we also conduct evaluations with human expert and non-expert raters and LLMs and study similarities and differences between these different sources of judgment.
@@
While most works have focused on evaluating model-generated stories on long-term coherence, relevance to premise, repetitiveness, and overall interestingness (?), recent studies have also evaluated the creativity of AI models in producing stories (??????). Chakrabarty et al. (?) generates short stories using LLMs based on plots from popular fictional works featured in the New York Times and performs a detailed expert evaluation of both the model-generated and original stories. Their findings reveal that LLMs fall considerably short of experienced writers in creating truly creative content. Tian et al. (?) similarly finds that LLM-generated stories are non-diverse and typically lack suspense and tension. However, these works largely focus on comparing models to award-winning professional writers while our work centers around comparing the creative story-writing abilities of models to average human writers.
Past work closest to ours is the work of Orwig et al. (?) which also evaluates creative short story generation in both average humans and LLMs using the same five-story generation task. However, our work differs in several major aspects. First, Orwig et al. (?) compares a collection of humans to only a single model (either GPT-3 or GPT-4) where model story variation is achieved by varying temperature values. This setup makes an implicit assumption of treating the same model with a different decoding parameter as equal to an individual human. However, it remains unclear whether the same findings will hold if a population of different models is compared to a population of humans. Therefore, our study focuses on evaluating collections of both different humans and LLMs. Second, Orwig et al. (?) scores creativity with an overall rating and link content to different human memory processes. In our study, we conceptualize creativity as a multifaceted concept and characterize it by the dimensions of novelty, surprise and value (?). Finally, Orwig et al. (?) collects creativity ratings from non-expert raters and GPT-4 while our work considers ratings from both non-expert and expert human raters as well as three LLM judges. We further study where differences between these three types of judges in ratings come from, by predicting creativity ratings from automated metrics across multiple dimensions.
@@ Methods
Write a creative short story using a maximum of five sentences. The story must include the following three words: {items}. However, the story should not be about {boring_theme}.
In the instructions above, items refer to the cue words and boring_theme corresponds to a typical or uncreative storyline that would first come to mind about those cue words. For example, a typical storyline for cue words stamp, letter, send could be stamping a letter and sending it. We include these hints in the instructions to increase the creativity of both human and LLM generated stories.
Human data were collected from $60$ participants ($43\%$ female, age: $M=38.8,SD=13.6$ years) on Prolific [^1] https://www.prolific.com/, a crowd-sourcing platform. Participants not adhering to instructions were removed, resulting in a total of $59$ participants.
For a fair comparison, model data were also collected from $60$ different models that are diverse in model size, training data and model architecture. The full list of models can be found in Appendix Tables 5 and 6. All models were prompted in zero-shot setting with a highly creative decoding setup ($temperature=0.7,top\_p=0.95$).
In total, we collected $480$ stories ($60$ stories for humans and models each across $4$ item sets). To make sure all stories are meaningful and comparable in length, we performed some preprocessing to remove outlier stories that have less than $3$ or more than $7$ sentences in them. This filtering step resulted in a total of $431$ stories for final evaluation. Final data statistics can be found in Appendix Table 10.
@@
We evaluate both human and model stories using various automated metrics that correspond to different dimensions of creativity. These measures are either common methods relying on the basic linguistic structure of sentences (e.g. n-grams, dependency trees) or metrics based on the notion of semantic distance that has been shown as an effective automated metric to evaluate creativity (??????). More specifically, we consider the following metrics:
@@
One of the major dimensions of creativity is the novelty aspect (?). It is typically defined as the measure of how different an artifact is from other known artifacts in its class (?). To compute the novelty of a story, we employ the novelty metric from Karampiperis et al. (?) which defines it as the average semantic distance between the dominant terms (i.e. lemmatized content words) of the story, compared to the average semantic distance of the dominant terms in all stories. More formally, let $S_{n}$ be a given story, $S_{G}$ a corpus of all stories and $T_{n}$ and $T_{G}$ set of dominant terms respectively for $S_{n}$ and $S_{G}$. Then similar to Johnson et al. (?), we can define the average semantic distance between the dominant terms for $S_{n}$ as follows:
@@
where $|F|$ refers to the number of fragments and $F_{i}$ is the $i$-th fragment. We employ this metric to compute a surprise metric for each generated story.
@@
We had two trained research assistants score each of the $431$ valid stories on creativity, originality, surprise and effectiveness. Inspired by the original Turing test (?), they are also asked to judge each story on whether it was created by human or an AI. The inter-rater reliability of their judgments ranged from good to excellent (ICC=.$62-.90$). Therefore, for each variable we compute composite scores by taking the means of the two human expert judges.
@@
We also conducted the same evaluation with an independent sample of $96$ non-expert judges recruited via Prolific ($49\%$ female, age: $M=39.8,SD=12.4$ years). To make sure no story evaluation result is dominated by an individual rater judgment, we collected 5 ratings for each story. However, due to the time and cost constraints, we performed this evaluation study on a subset of the stories ($n=273$) that nonetheless cover all item sets. The variation in judgments across non-experts was higher than between experts or LLMs, where the inter-rater reliability of their judgments ranged from fair to good (ICC=.$43-.71$). Therefore, we choose to use the median rating (i.e., most common rating given across all ratings of the story) rather than a mean to reduce the influence of outliers.
@@ Results
Moreover, inverse homogenization scores indicate that model stories tend to share the same themes while human stories are much more diverse in their content. To quantify the diversity of themes, similar to Nath et al. (?) we perform agglomerative clustering using Ward variance minimization algorithm with a threshold of $0.6$ on the embeddings of all stories for a given item set which gives us a set of theme clusters. We find that human stories are characterized by significantly more themes than model stories (2x more on 3 out of 4 item sets). To further analyze what type of themes dominate model and human stories, for each group, we combine stories written on a given item set and ask GPT-4 to summarize them in a sentence. Results show that model stories tend to focus on themes of magical transformations or mysterious transactions while human stories speak of human nature, human interactions and responsibilities. Theme statistics and identified categories can be found in Appendix Figure 14 and Table 4 respectively.
@@
Figures 5 and 6 summarize some of the results for lexical and syntactic complexity metrics respectively across all model and human groups and item sets. What we see is that AI models consistently produce longer stories and their sentences are syntactically more complex as indicated by larger number of unique words and longer dependency paths per sentence ($p<0.001$). Additionally, we find that models generally use more nouns and adjectives, while humans use more pronouns and adverbs ($p<0.001$). To further analyze the type of pronouns used by humans and AI models, we perform an additional analysis on pronoun use and find that humans almost exclusively write their stories from the first or second person perspective, however, models prefer stories centered around third person. Overall, our findings show that models generally produce grammatically complex and potentially less readable stories. Results for pronoun analysis are in Appendix Figure 11. More complexity metric results are in Appendix Figures 8 and 9.
@@
When we analyze the effect of the item set semantic distance on the complexity scores, we observe a significant difference only with respect to lexical complexity scores for human stories where low semantic distance results are more lexically complex than high semantic distance results ($p<0.01$). Results for this analysis are in Appendix Figure 10.
@@
Experts predict the author (human or AI) correctly $94\%$ of the time, outperforming both non-expert and LLM judges. Non-experts make the correct prediction $81\%$ of the time and LLM judges predict human vs. AI with $71\%$ accuracy. Additionally, to gain insight into what factors drive human judgment on whether a story is written by AI or a human, we ask the non-experts to explain their strategy to predict the author of a story. We summarize their comments into several themes of qualities that were attributed the most respectively to human and AI stories. We find that qualities identified by non-experts coincide with our findings from the automated metric analysis that AI models tend to produce linguistically more complex and verbose yet less creative stories than humans. Qualities identified by non-experts can be found in Appendix Table 9.
@@ Discussion
Our analysis using the automated metrics shows that LLMs produce linguistically and stylistically more complex stories than humans as indicated by higher lexical and syntactic complexity results. However, human stories consistently exhibit higher novelty, surprise, and lexical and semantic diversity while being linguistically much less complex and easier to read. Our findings are in line with some previous work comparing LLM creativity to humans in story generation (????) and creative problem-solving (?). On the other hand, some past works have found no significant difference between overall LLM and human creativity in short story generation when comparing a population of humans to GPT-3 and GPT-4 when judged by non-experts and GPT-4 (?). However, our fine-grained analysis considering multiple dimensions of creativity evaluated by automated metrics and a population of different LLMs reveals significant gaps between humans and LLMs across all major dimensions of creativity in favor of humans.
@@
Additionally, we find that our automated metric results highly correlate with expert judgments, while LLM and non-expert judgments tend to rate LLM stories as more creative than humans. However, our further analysis shows that this discrepancy stems from the underlying factors driving the different judgments. More specifically, expert judgments are driven by the diversity and surprise aspects of the stories which are essential to creativity while non-expert and LLM judgments highly correlate with sheer lexical complexity such as the number of unique words, which does not necessarily imply semantic complexity. In fact, our parts-of-speech analysis shows that LLMs tend to overuse rare adjectives and complex syntactical structures. Additionally, separating the difficulty of technical skill from the creativity of the content when judging creativity is a also problem in artistic creativity (?). Moreover, past works generally find expert judgments as more reliable evaluators of creativity than those of non-experts (???????). Similarly, LLMs-as-judges have been shown to be unreliable (??) and biased towards their own generations (??).
@@
Our work has several implications. The complexity vs. creativity gap shows that humans and LLMs have different understandings of what it means to be creative for stories. Examples of human and AI stories that are deemed the most and least creative by experts can be found in Appendix Tables 7 and 8 respectively. While humans prefer telling a simple story from their perspective that is nonetheless surprising and original, LLMs however, confuse creativity with lexically and syntactically overloaded sentences narrated from the third person perspective and that typically focus on a few repetitive themes. This behaviour can be due to several factors involved in training LLMs such as the training data, pre-training and post-training optimizations. For example, aligning LLMs with human feedback to be more helpful has been attributed to result in strong verbosity bias (?) and diversity reduction (?) in creative tasks. This bias also pushes LLMs to prefer linguistically complex responses as a judge, which could be problematic for automated metrics using LLMs (?). Our findings call for a more comprehensive evaluation of creativity and can inform future work on designing methods to improve the creativity of LLMs (?). Potential directions can include developing new prompt engineering (????) or optimization techniques (???) or steering internal mechanisms of LLMs using mechanistic interpretability (?).

### v5

@@ Abstract
Story-writing is a fundamental aspect of human imagination, relying heavily on creativity to produce narratives that are novel, effective, and surprising. While large language models (LLMs) have demonstrated the ability to generate high-quality stories, their creative story-writing capabilities remain under-explored. In this work, we conduct a systematic analysis of creativity in short story generation across 60 LLMs and 60 people using a five-sentence cue-word-based creative story-writing task. We use measures to automatically evaluate model- and human-generated stories across several dimensions of creativity, including novelty, surprise, diversity, and linguistic complexity. We also collect creativity ratings and Turing Test classifications from non-expert and expert human raters and LLMs. Automated metrics show that LLMs generate stylistically complex stories, but tend to fall short in terms of novelty, surprise and diversity when compared to average human writers. Expert ratings generally coincide with automated metrics. However, LLMs and non-experts rate LLM stories to be more creative than human-generated stories. We discuss why and how these differences in ratings occur, and their implications for both human and artificial creativity.
@@ Introduction
Story-writing lies at the core of human imagination and communication, serving as a potent means to connect and convey ideas effectively spanning across all human cultures and time periods (?). It typically demands creativity, especially when shaping a captivating and persuasive narrative. Creativity is the ability to produce novel, useful, and surprising ideas, and has been widely studied as a crucial aspect of human cognition (????). While humans are natural storytellers, getting machines to generate stories automatically has been a long-time challenge (?????). However, recently large language models (?) have been shown to produce high-quality short and long stories on arbitrary topics (??). These stories are often evaluated by humans on their global coherence, relevance to the premise, repetitiveness, and general interestingness to the reader (?). However, the extent to which these model-generated stories are truly creative, i.e. novel, effective, and surprising, remains under-studied.
@@
Recent works evaluating LLM’s ability to produce creative content have shown that models largely fall behind professional human writers (????). On the other hand, Orwig et al. (?) finds no significant difference between average human and AI-generated short stories in terms of creativity ratings by non-experts or GPT-4, when comparing humans to ChatGPT models. The extent to which these results hold when considering a collection of different models and evaluations across multiple dimensions of creativity as well as expert and non-expert raters remains unclear.
To bridge these gaps, in this work, we conduct a systematic analysis of creativity in short story generation in humans and LLMs. We employ a creative short story generation task that is typically used in psychology to measure the creativity of humans (???). In this task, the goal is to write a short creative story in approximately five sentences based on three cue words such as stamp, letter and send. We evaluate 60 humans and 60 state-of-the-art instruction-finetuned large language models on this task and analyze their performance based on multiple automatic metrics of creativity representing, overall creativity, and the specific aspects of novelty, surprise, diversity, and complexity. Our analysis shows that model-generated stories tend to employ more complex linguistic structures than humans; however, they significantly fall short when it comes to novelty, diversity, and surprise compared to average human writers.
@@
Evaluating creativity is a challenging task due to its subjective nature, however, several evaluation methods have been proposed in the past (??). The most common method of creativity evaluation is called the Consensual Assessment Technique (CAT) (?). CAT relies on the collective judgment of human experts. However, the level of expertise required for a human rater is a subject of debate, with most evidence favoring (quasi-)experts over non-experts as good raters (???????). With the rise of powerful generative AI models, LLMs are increasingly being used as judges in many evaluation tasks including creativity (??????). Consequently, in addition to automated metrics, we also conduct evaluations with human expert and non-expert raters and LLMs and study similarities and differences between these different sources of judgment.
@@
While most works have focused on evaluating model-generated stories on global coherence, relevance to premise, repetitiveness, and overall interestingness (?), recent studies have also evaluated the creativity of AI models in producing stories (??????). Chakrabarty et al. (?) generates short stories using LLMs based on plots from popular fictional works featured in the New York Times and performs a detailed expert evaluation of both the model-generated and original stories. Their findings reveal that LLMs fall considerably short of experienced writers in creating truly creative content. Tian et al. (?) similarly finds that LLM-generated stories are non-diverse and typically lack suspense and tension. However, these works largely focus on comparing models to award-winning professional writers while our work centers around comparing the creative story-writing abilities of models to average human writers.
Past work closest to ours is the work of Orwig et al. (?) which also evaluates creative short story generation in both average humans and LLMs using the same five-story generation task. However, our work differs in several major aspects. First, Orwig et al. (?) compares a collection of humans to only a single model (either GPT-3 or GPT-4) where model story variation is achieved by varying temperature values. This setup makes an implicit assumption of treating the same model with a different decoding parameter as equal to an individual human. However, it remains unclear whether the same findings will hold if a population of different models is compared to a population of humans. Therefore, our study focuses on evaluating collections of both different humans and 60 different LLMs. Second, Orwig et al. (?) scores creativity with an overall rating and links content to different human memory processes. In our study, we conceptualize creativity as a multifaceted concept and characterize it by the dimensions of novelty, surprise and value (?). Finally, Orwig et al. (?) collects creativity ratings from non-expert raters and GPT-4 while our work considers ratings from both non-expert and expert human raters as well as three LLM judges. We further study where differences between these three types of judges in ratings come from, by predicting creativity ratings from automated metrics across multiple dimensions.
@@ Methods
Write a creative short story using a maximum of five sentences. The story must include the following three words: {items}. However, the story should not be about {boring_storyline}.
In the instructions above, items refer to the cue words and boring_storyline corresponds to a typical or uncreative storyline that would first come to mind about those cue words. For example, a typical storyline for cue words stamp, letter, send could be stamping a letter and sending it. We include these hints in the instructions to increase the creativity of both human and LLM generated stories.
Human data were collected from $60$ participants ($43\%$ female, age: $M=38.8,SD=13.6$ years; fluent English speakers residing in the UK, with no language-related disorders and having completed secondary school education) on Prolific [^1] https://www.prolific.com/, a crowd-sourcing platform. Participants not adhering to instructions were removed, resulting in a total of $59$ participants.
For a fair comparison, model data were also collected from $60$ different models that are diverse in model size, training data and model architecture. The full list of models can be found in Models section. All models were prompted in zero-shot setting with a decoding setup that has been used in previous works to generate creative outputs ($temperature=0.7,top\_p=0.95$) (??).
In total, we collected $480$ stories ($60$ stories for humans and models each across $4$ item sets). To make sure all stories are meaningful and comparable in length, we performed some preprocessing to remove outlier stories that contain less than $3$ or more than $7$ sentences. This filtering step resulted in a total of $431$ stories for final evaluation. Final data statistics can be found in Appendix Table 10.
@@
We evaluate both human and model stories using various automated metrics that correspond to different dimensions of creativity. These measures are either common methods relying on the basic linguistic structure of sentences (e.g. n-grams, dependency trees) or metrics based on the notion of semantic distance that has been shown as an effective automated metric to evaluate creativity (??????). Semantic distance between two texts is typically computed based on the cosine similarity of embeddings of the texts. More specifically, we consider the following metrics:
@@
One of the major dimensions of creativity is the novelty aspect (?). It is typically defined as the measure of how different an artifact is from other known artifacts in its class (?). To compute the novelty of a story, we employ the novelty metric from Karampiperis et al. (?) which defines it as the average semantic distance between the dominant terms (i.e. lemmatized content words) of the story, compared to the average semantic distance of the dominant terms in all stories. More formally, let $S_{n}$ be a given story, $S_{G}$ a corpus of all stories across all item-sets and $T_{n}$ and $T_{G}$ set of dominant terms respectively for $S_{n}$ and $S_{G}$. Then similar to Johnson et al. (?), we can define the average semantic distance between the dominant terms for $S_{n}$ as follows:
@@
where $|F|$ refers to the number of fragments and $F_{i}$ is the $i$-th fragment. We employ this metric to compute a value of surprise for each generated story.
@@
We had two trained research assistants—graduate psychology students with creative writing experience— score each of the $431$ valid stories on creativity, originality, surprise and effectiveness. For unbiased evaluation, these annotators were not involved in any stage of the study, were unfamiliar with the study design and were fully blinded to which stories stemmed from humans versus AI. Inspired by the original Turing test (?), they were also asked to judge each story on whether it was created by human or an AI. The inter-rater reliability of their judgments ranged from good to excellent (ICC=.$62-.90$). Therefore, for each variable we compute composite scores by taking the means of the two human expert judges.
@@
We also conducted the same evaluation with an independent sample of $96$ non-expert judges recruited via Prolific ($49\%$ female, age: $M=39.8,SD=12.4$ years) The non-expert raters were -just like the recruited amateur writers- UK residents who spoke English fluently, had no language-related disorders and have completed secondary education. We collected 5 non-expert ratings for each story, the minimum required for reliable creativity judgments by non-experts (?). Due to time and cost constraints, we performed this evaluation study on a subset of the stories ($n=273$) that nonetheless cover all item sets.
To ensure consistent and reliable evaluation, all annotators received detailed instructions on the definitions of creativity, novelty, surprise, and effectiveness. As often is the case in subjective creativity judgments, the variation in judgments across non-experts was much higher than between experts or LLMs (?), where the inter-rater reliability of their judgments ranged from fair to good (ICC=.$43-.71$). Therefore, we choose to use the median rating (i.e., most common rating given across all ratings of the story) rather than a mean to reduce the influence of outliers.
@@ Results
Moreover, inverse homogenization scores indicate that model stories tend to share the same themes while human stories are much more diverse in their content. To quantify the diversity of themes, similar to Nath et al. (?) we perform agglomerative clustering using Ward variance minimization algorithm with a threshold of $0.6$ on the embeddings of all stories for a given item set which gives us a set of theme clusters. We find that human stories are characterized by significantly more themes than model stories (Figure 8). To further analyze what type of themes dominate model and human stories, for each group, we combine stories written on a given item set and ask GPT-4 to summarize them in a sentence. Results show that model stories tend to focus on themes of magical transformations or mysterious transactions while human stories speak of human nature, human interactions and responsibilities. Theme statistics and identified categories can be found in Appendix Figure 8 and Table 5 respectively.
@@
Figures 5 and 6 summarize some of the results for lexical and syntactic complexity metrics respectively across all model and human groups and item sets. What we see is that AI models consistently produce longer stories and their sentences are lexically and syntactically more complex as indicated by larger number of unique words and longer dependency paths per sentence ($p<0.001$). Additionally, we find that models generally use more nouns and adjectives, while humans use more pronouns and adverbs ($p<0.001$). To further analyze the type of pronouns used by humans and AI models, we perform an additional analysis on pronoun use and find that humans almost exclusively write their stories from the first or second person perspective, however, models prefer stories centered around third person. Overall, our findings show that models generally produce grammatically complex and potentially less readable stories. Results for pronoun analysis are in Appendix Figure 12. More complexity metric results are in Appendix Figures 9 and 10.
@@
When we analyze the effect of the item set semantic distance on the complexity scores, we observe a significant difference only with respect to lexical complexity scores for human stories where low semantic distance item sets result in more lexically complex stories than high semantic distance item sets ($p<0.01$). Results for this analysis are in Appendix Figure 11.
@@
Experts predict the author (human or AI) correctly $94\%$ of the time, outperforming both non-expert and LLM judges. Non-experts make the correct prediction $81\%$ of the time and LLM judges predict human vs. AI with $71\%$ accuracy. Additionally, to gain insight into what factors drive human judgment on whether a story is written by AI or a human, we ask the non-experts to explain their strategy to predict the author of a story. We summarize their comments into several themes of qualities that were attributed the most respectively to human and AI stories. We find that qualities identified by non-experts coincide with our findings from the automated metric analysis that AI models tend to produce linguistically more complex and verbose yet less creative stories than humans (Table 4).
@@ Discussion
Our analysis using the automated metrics shows that LLMs produce linguistically and stylistically more complex stories than humans as indicated by higher lexical and syntactic complexity results. However, human stories consistently exhibit higher novelty, surprise, and lexical and semantic diversity while being linguistically much less complex and easier to read. Our findings are in line with some previous work comparing LLM creativity to humans in story generation (????) and creative problem-solving (?). On the other hand, some past works have found no significant difference between overall LLM and human creativity in short story generation when comparing a population of humans to GPT-3 and GPT-4 when judged by non-experts and GPT-4 (?). However, our fine-grained analysis considering multiple dimensions of creativity and a population of 60 different LLMs evaluated by automated metrics and experts reveals significant gaps between human and LLM stories across all major dimensions of creativity in favor of humans.
@@
Particularly, we find that our automated metric results highly correlate with expert judgments, while LLM and non-expert judgments tend to rate LLM stories as more creative than human stories. Our further analysis shows that this discrepancy stems from the underlying factors driving the different judgments. More specifically, expert judgments are driven by the diversity and surprise aspects of the stories which are essential to creativity while non-expert and LLM judgments highly correlate with sheer lexical complexity such as the number of unique words, which does not necessarily imply semantic complexity. In fact, our parts-of-speech analysis shows that LLMs tend to overuse rare adjectives and complex syntactical structures. Moreover, past works generally find expert judgments as more reliable evaluators of creativity than those of non-experts (???????). Similarly, LLMs-as-judges have been shown to be unreliable (??) and biased towards their own generations (??).
Figure 8: Number of themes for human and AI stories.
@@
Our work has several implications. The complexity vs. creativity gap shows that humans and LLMs have different interpretations of what it means to be creative for stories. While humans prefer telling a simple story from their perspective that is nonetheless surprising and original, LLMs, however, represent creativity with lexically and syntactically overloaded sentences narrated from the third person perspective and that typically focus on a few repetitive themes. Additionally, the fact that non-experts and LLMs tend to evaluate AI stories as more creative could mean that complexity creates the illusion of being more creative to the untrained eye. This behaviour can be due to several factors involved in training LLMs such as the training data, pre-training and post-training optimizations. For example, aligning LLMs with human feedback to be more helpful has been attributed to result in strong verbosity bias (?) and diversity reduction (?) in creative tasks. Our findings call for a more comprehensive evaluation of creativity and can inform future work on designing methods to improve the creativity of LLMs (?). Potential directions can include developing new prompt engineering (????) or optimization techniques (???) or steering internal mechanisms of LLMs using mechanistic interpretability (?).
Limitations
While our study provides a comprehensive analysis of a wide range of language models using fine-grained creativity metrics, it does have a few limitations. First, although all models were prompted using a decoding setup that has been used in previous work to favor creative output and idea diversity (??), we did not explore alternative decoding or prompting strategies due to the high cost of open-ended evaluation across many models. Second, the novelty metric we employ is reference-based, meaning its outcomes depend heavily on the chosen reference corpus. In our case, this was the set of human- and AI-written stories for each item set. Using reference-based scoring is a shortcoming in nearly all creativity research that uses uniqueness to define novelty (?). However, the construct validity of frequency-based uniqueness assessments is also high and their generalizability to stories from other populations increases with sample size (?). Therefore, future studies could validate our findings by including more stories in reference-based metrics. Third, we used the minimum recommended number of expert and non-expert raters per story, i.e., two for (quasi-)experts —as these are generally highly reliable as in our study— and five non-experts —as these are generally less reliable— (?). Having more raters could have improved the reliability of our findings. Furthermore, including professional creative writers as experts could perhaps provide further improve reliability (although research suggests that quasi-experts are as reliable as experts (?), sometimes more so (?)). Lastly, although our creativity metrics are effective for evaluating story generation, they cannot fully capture the broader cultural and social dimensions of creativity or the depth of truly original and imaginative language use.